Suppose the Swiss Franc exchange rate is CHF 2.00 = USD 1.00 and the British Pound exchange rate is GBP 0.60 = USD 1.00.

Suppose a dealer quotes a rate of 3 CHF = 1 GBP.

  1. What is the CHF/GBP cross rate?
  2. Is there an arbitrage opportunity here? If, yes then calculate the arbitrage profit on $5,000?
Answer #1

CHF/GBP Cross Rate = (CHF/USD)*(USD/GBP)

= 2*(1/0.6)

= CHF 3.33333/GBP

Since actual rate is different from cross rate, arbitrage opportunity exists

Convert USD into CHF and get 5,000*2 = 10,000 CHF

Convert into GBP and get 10,000/3 = 3,333.33

Convert back into USD and get 3,333.33/0.6 = $5,555.56

Hence, arbitrage profit = $555.56